For what value of x; x^4+4x^2+x+8/x^2+4/x^4 is a complete square where x is not zero
1
Expert's answer
2020-06-21T17:23:55-0400

x4+4x2+x+8/x2+4/x2=x4+4/x4+4-4+x+4x2+8/x2

as

x4+4/x4+4=(x2+2/x2)2 : 4x2+8/x2=4(x2+2/x2)

we get

(x2+2/x2)2+4(x2+2/x2)-4+x

complete square

(x2+2/x2)2+4(x2+2/x2) +4

where from 4=-4+x

x=8